Please solve and show answers so I can do it on my own
Use the following information to answer questions 15 through 24 The systolic blood pressure of individuals is thought to be related to both age and weight. A random sample of 11 men was taken. The Data and the Excel computer output for the multiple regression problem is given below Pressure Age Weight 132 143 153 162 154 168 137 149 159 128 166 52 59 67 73 64 74 54 61 65 46 72 173 184 194 211 196 220 188 188 207 167 217 SUMMARY OUTPUT Regression Statistics Multiple R 0.988356 R Square 0.976847 Adjusted 0.971059 R Square Standard 2.318211 Error Observat
Residual 8 42.93282 5.374103 906.9581 188 7646 28TE-07 Pyalue Lower Upper Lower 0.031865 3.451678 58.53653 3.451678 58.53853 Upper 95 Standard Stat Intercept30.9941 11 X Variable 0.861415 0.248231 X Variable 0334850 01 1433838 11.94378 2.504909 861415 0.248231 3.470208 0 .00844 0.288992 0.00844 0.288092 1.433838 0280992 15) Waat is the estimated 16) What is the value of test statit relationahip amongthe vaziablea if there iaa used to determine 17) The cri is a relationship among the vaxiabies a) 2.96 b) 3.59 c) 3.16 d) 2.85 e) 4-10 E) 4.76 g) 4.46 tical F value obtained from the table which in used to test if theze table at the s level equalai 18) Carry out the test to determine if there is a relationship among the variables at 5s level. The null hypothesis should be a) rejected b) not rejected c) impossible to determine d) none of the above answers aze correct 20) Given x,-45, x-200, use the estimated regression equation to estimate the systolic blood pressure. si8. 19) what is the null hypothesia, or for problem 16) 21) We,want to test vhether the parameter B is significant. What does the test statistioc equal?
the positive t value obtained from the table which is used to test an individual parameter at the 5% level is: a) 1.753 b) 1.756 c) 2.602 d) 2.583 E) 2.120 h) none of the above ansvers are correct B at the 5t level. The nul 23) Carry out the test of significance for the parameter hypothesis should be: a) rejected b) not rejected c) impossible to determine d) none of the above answers are correct 24) The null hypothesis, Ho, for problem 21
Answer #1

Let me answer some of the questions for you-

15) Based on the output, the regression output can be given as follows-

Y = 30.99 + 0.961*X1 + 0.334*X2

16) For multiple regression, the test statistic used is: F-value

17) F Critical at 5% level of significance for 10 degrees of freedom is given from the F-table as

F Critical = 2.98 which will be (a)

19)

Ho: Null Hypothesis will be: b1=b2=0

20) X1 = 45 and X2 = 200

Y = 30.99 + 0.961*45 + 0.334*200 = 141.03

Let me know if you need anything else, if not please don't forget to like the answer :)